A red dy 2 m length. and a Square (2.o mn x 2, o mm) CT ss section is made d al wish a resistivity ds 6xlo J2 m. I a potential difference 0.50 v is pla ted across she ends dd ne rad, a what rate is heat ene rated in rod 2
Answer #1

resistance of rod will be given as :

R = rho L / A

R = (6 x 10^-4 ohm m ) (2 m) / (2 x 10^-3 x 2 x 10^-3 m^2)

R = 300 ohm


P = V^2 / R = 0.50^2 / 300


P = 8.33 x 10^-4 W Or 0.833 mW